OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (1, 2)?

OpenStudy (jhannybean):

The vertex form of a parabolic equation follows the format \[\large y= \color{purple}a(x-\color{green}h)^2 +\color{green}k\]

OpenStudy (anonymous):

??

OpenStudy (jhannybean):

You havent finished your questions....there are no solutions to choose from?

OpenStudy (anonymous):

y = (x - 1)2- 2 y = (x - 1)2+ 2 y = (x + 1)2- 2 y = (x + 1)2+ 2

OpenStudy (jhannybean):

Ok, in this case our \(a = 1\), so we don't have to worry about that. :) The coefficient in front of \((x-h)^2\), our "a" value,tells us which way the parabola opens, in this case, all our answer choices indicate its opening upwards. Now we're given points of our VERTEX which is represented by \(\color{red}{(h,k)}\) In thi case, our vertex = \(\color{green}{(1,2)}\)

OpenStudy (jhannybean):

If we follow the format of the parabolic equation in vertex form from the first post, we see that \(\large y = (x-\color{green}h)^2 +\color{green}k \iff y= (x-\color{green}1)^2 +\color{green}2\)

OpenStudy (anonymous):

im confused lol

OpenStudy (anonymous):

????

OpenStudy (jhannybean):

With which part?

OpenStudy (jhannybean):

Ok, the standard form of a parabola is \(y=ax^2+bx+c\).You understand this, right?